OpenStudy (anonymous):

a bag contains 10 white golf balls and 6 striped golf balls. a golfer want to add 112 golf balls to the bag. he wants the ratio to stay the same. how many of each should he add?

OpenStudy (anonymous):

someone please help

OpenStudy (natsent06):

ratio white : striped = 10 : 6 10x+6x = 16+112 = 128 16x = 128 x = 8 so there are now 80 white and 48 striped white to add = 80-10 = 70 .....answer striped to add = 48-6 = 42 ......answer

OpenStudy (anonymous):

thank you